A ball is thrown downward from the top of a 90 story building. The distance, d inf feet, the ball is above the ground after t secconds is given by the function equation d=f(t) = -16t^2 -80t +800. After 2 seconds whaat is the distance of the ball from the ground? What is the height of the building, in feet? How long does it take for the ball to hit the ground?
I do know that the distance from the ground is 576 feet. That is if it is correct.

Answer by lwsshak3(11628)

A ball is thrown downward from the top of a 90 story building. The distance, d inf feet, the ball is above the ground after t seconds is given by the function equation d=f(t) = -16t^2 -80t +800. After 2 seconds whaat is the distance of the ball from the ground? What is the height of the building, in feet? How long does it take for the ball to hit the ground?
I do know that the distance from the ground is 576 feet. That is if it is correct.
**
d=f(t) = -16t^2 -80t +800
after 2 seconds, height of ball
f(2)=-16(2^2)-80(2)+800
=-64-160+800=576 feet
..
Height of Bldg
At 0 seconds the ball has dropped 0 feet from the top of the building.
Height of building=f(0)=800 feet
..
When ball hits the ground, its distance above the ground=0
0=-16t^2 -80t +800
-16t^2 -80t +800=0
t^2+5t-50=0
(t+10)(t-5)=0
t=-10 (reject, t>0)
t=5 seconds
ans:
distance of the ball from the ground after 2 seconds=576 feet
Height of building=800 feet
How long it takes for the ball to hit the ground=5 seconds